OpenStudy (anonymous):

Find the unit rate in feet per second. 180miles/2 hours

OpenStudy (anonymous):

180 miles = 180 (feet per mile) 2 hours = 2 (seconds in an hour)

OpenStudy (anonymous):

figure those out and divide

OpenStudy (amistre64):

\[\frac{180\ miles}{2\ hours}*\frac{1\ hour}{60\ min}*\frac{1\ min}{60\ sec}*\frac{5280\ feet}{1\ mile}\] \[\frac{180\ \cancel{miles}}{2\ \cancel{hours}}*\frac{1\ \cancel{hour}}{60\ \cancel{min}}*\frac{1\ \cancel{min}}{60\ sec}*\frac{5280\ feet}{1\ \cancel{mile}}\] \[\frac{18\cancel{0}*528\cancel{0}\ feet}{2*6\cancel{0}*6\cancel{0}\ sec}\] \[\frac{18*528\ feet}{2*6*6\ sec}\] etc...

OpenStudy (anonymous):

\[\frac{90 \text{ miles}}{1 \text{ hour}}=\frac{90*5280 \text{ feet}}{1*3600 \text{ seconds}}=132 \text{ ft}/\sec \]